Find the relative extrema for


i) 𝑓( π‘₯ )= π‘₯^3 βˆ’ 3π‘₯ + 5


𝑖𝑖) f(x)=π‘₯^4 + 2π‘₯^2 βˆ’ 4

1
Expert's answer
2022-03-25T06:33:16-0400

i) "f(x)=x^3-3x+5"

"f\u2019(x)=3x^2-3=0"

"x=\pm1" ;

"f\u2019(-2)=12-3>0"

"f\u2019(0)=-3<0"

"f\u2019(2)=12-3>0"

Thus "x=-1" - maximum; "x=1" - minimum.

Answer: "x_{max}=-1" ; "x_{min}=1" .

ii) "f(x)=x^4+2x^2-4"

"f\u2019(x)=4x^3+4x=4x(x^2+1)=0"

"x=0"

"f\u2019(-1)=-8<0"

"f\u2019(1)=8>0"

"x=0" - minimum.

Answer: "x_{min}=0" .